100 cm³ of 2.0 mol dm‐³ hydrochloric acid is poured into 100 cm³ of 2.0 mol dm‐³ sodium hydroxide solution. The initial and final temperature are immediately recorded as shown below. [Specific heat capacity = 4.2 J/g°C, density of solution = 1.0 gcm-3]. What is the value of heat released in the reaction?

initital temperature of hydrochloric acid: 29.0°C

initial temperature of sodium hydroxide solution: 29.0°C

highest temperature of the mixture: 35.0°C
1
Expert's answer
2020-05-20T13:46:52-0400

1) HCl + NaOH → NaCl + H2O 

Moles (HCl) = Conc x Vol/1000 = 2.0*100/1000 = 0.2 moles

Moles (NaOH) = Conc x Vol/1000 = 2 * 100/1000 = 0.2 moles

2) Q = m*c*ΔT

Total mass of "water" = 200 g

ΔT = 35 - 29 = 6 K

Q= 200*6*4.2=5.04 kJ

3) ΔHneutralisation =Q/moles of water made

ΔHneutralisation=5.04/0.2 = 25.2 kJ